### v1

@@ 2.1 Classic Proper Orthogonal Decomposition (POD)
The classical Proper Orthogonal Decomposition (POD) is a widely used technique for extracting low-dimensional structures from high-dimensional spatiotemporal data [berkooz1993proper, chatterjee2000introduction, volkwein2013proper, kunisch2008proper]. It is especially prominent in fluid dynamics, model reduction, and system identification due to its mathematical rigor, computational efficiency, and optimality properties in the $L^{2}$ norm.
@@
subject to the orthonormality constraint $(\phi_{i},\phi_{j})=\delta_{ij}$. The solution is obtained via singular value decomposition (SVD) of the snapshot matrix, and the resulting modes $\phi_{i}$} span an $r$-dimensional subspace that captures the most energetic components of the dataset [gubisch2017proper, kunisch2008proper, berkooz1992observations]. POD guarantees optimality in the $L^{2}$ sense: for any given rank $r$, the POD basis minimizes the projection error among all $r$-dimensional linear subspaces. This makes it a powerful tool for linear dimensionality reduction and the foundation for many reduced order models (ROMs), where the governing equations are projected onto the POD subspace using Galerkin or Petrov–Galerkin projection [girfoglio2021pod, besabe2025data, iliescu2013variational, mou2020data, mou2020data, tsai2025time].
Despite its strengths, classical POD has several limitations. Because its basis is strictly linear, it often struggles to represent complex nonlinear features, sharp gradients, or discontinuities [li2016performance, volkwein2013proper, iliescu2014are]. Furthermore, the POD basis is highly dependent on the resolution and parameter regime of the snapshot data, limiting its generalizability [mifsud2016variable, roderick2014proper]. These limitations (see left panel of Figure 1(a)) motivate the development of more flexible, data-driven alternatives, such as the Neural-POD framework introduced in this work (see right panel of Figure 1(a)).
2.2 Neural-POD Framework
@@ 2.2.2 Loss Functions
Figure 2: Overview of Neural-POD training. Snapshots $u(x,t,\kappa)$ are collected over time for varying parameter values $\kappa$ (upper left). The training architecture decomposes each snapshot into products of spatial mode functions $\Phi_{i}(x;\kappa)$ and temporal coefficients $\Psi_{i}(t;\kappa)$, which guarantees the orthogonality for all modes (bottom left). Once trained, the Neural-POD basis provides plug-and-play reduced representations that can be evaluated on any discretization $\hat{x}$ and parameter choice $\hat{\kappa}$ (right, $\hat{\kappa}$ may differ from the training snapshots’ parameter $\kappa$).
2.3 Neural-POD in Galerkin Projection Based Reduced Order Models
@@
Figure 3: Schematic diagram of the proposed Neural-POD-ROMs and traditional POD-based reduced order models (POD-ROMs). Top portion: the Neural-POD extension that enables plug-and-play operation on arbitrary meshes and at arbitrary resolution, informed by pretrained neural representations of POD bases and PDE parameters. Bottom portion: the traditional POD flowchart, including snapshot generation for parameter $\kappa$, computation of full POD modes via SVD, construction of the reduced basis, and Galerkin projection of the governing PDE.
2.3.1 Traditional Projection-ROM
@@ 2.3.2 Neural-POD-ROM
where the tensors $C_{ijk}$ and $D_{ik}$ are now computed using the Neural-POD modes. The projection-based ROM using POD and Neural-POD is specifically illustrated in Figure 3.
@@
2.4 Neural-POD in Deep Operator Networks
Figure 4: Schematic of the proposed Neural-POD Deep Operator Network (NP-DeepONet). The branch network takes initial condition fields as input, while the trunk network evaluates functions at spatial query points. A pretrained Neural-POD module provides reduced basis information $\phi_{1},\phi_{2},\ldots,\phi_{P}$} obtained from snapshot data. The outputs are combined to produce the operator evaluation $G_{\theta}(u_{m})(x)$, enabling efficient and data-driven surrogate modeling of PDE solutions in reduced coordinates.
@@ 2.4 Neural-POD in Deep Operator Networks
where $\phi_{k}(\cdot;\nu,\mathcal{X})$ denote the pretrained Neural-POD modes evaluated at viscosity $\nu$ and spatial grid $\mathcal{X}=\{x_{i}\}_{i=1}^{N}$, and $\phi_{0}$ is the corresponding mean field. The branch network outputs $b_{k}(v)$ represent the modal coefficients associated with these Neural-POD basis functions. The Neural-POD-DeepONet framework is illustrated in Figure 4.
3 Numerical Results
3.1 Burgers’ equation
$u(x,0)=u_{0}(x),$ $x\in(0,1),$ (19)
In the following experiments, the Neural-POD model is trained using snapshot data generated at viscosity values $\nu$ sampled within the interval $[0.005,,0.01]$. This parameter range defines the training domain for the Neural-POD basis learning. Figure 5 shows the evolution of the one-dimensional Burgers equation for different sampled viscosity values $\nu$ used to construct the training dataset for Neural-POD. As $\nu$ decreases (right to left), the solution transitions from a smooth diffusive profile to one dominated by advective nonlinearities, leading to the formation of sharp fronts and nearly discontinuous shock structures. This coupled effects of convection and diffusion underscores the limits of traditional POD in handling unseen parameters, motivating parametric POD formulations that adapt to viscosity and smoothness variations. More details can be found in Appendix 5.1.
Figure 5: Solution snapshots of the Burgers equation at different sampled viscosity values $\nu$ used for Neural-POD training.
3.1.1 Neural-POD with $L^{1}$ and $L^{2}$ loss
Figure 6 compares the first three spatial modes obtained from the classical POD and those learned through Neural-POD training using $L^{1}$- and $L^{2}$-based losses across different viscosity regimes. The $L^{2}$-optimal Neural-POD tends to emphasize smooth, energy-dominant structures, while the $L^{1}$-based formulation better preserves localized sharp transitions by penalizing large pointwise deviations less severely. Figure 7 shows that the Neural-POD residual decreases rapidly with the number of modes, closely mirroring the decay of the classical POD residual. More details are provided in Section 5.1.
Figure 6: First three traditional POD modes and Neural-POD under different loss choices, i.e., $L^{1}$ and $L^{2}$ norms, for different viscosity values.
Figure 7: Neural-POD residuals versus modes $r$ for the $L^{2}$ loss (left) and the $L^{1}$ loss (right).
Figure 9 compares the $L^{1}$ and $L^{2}$ norms of the reconstruction residuals obtained from Neural-POD models trained with $L^{1}$- and $L^{2}$-based loss functions across different viscosity values. Panels (a) and (b) correspond to the $L^{2}$ and $L^{1}$ residual norms, respectively, evaluated for three representative cases of $\nu=0.01$, $\nu=0.03$, and $\nu=0.05$. In panel (a), the $L^{2}$-optimal Neural-POD produces smooth and globally accurate reconstructions, with residuals remaining low throughout the spatial domain, particularly for higher viscosity values where the flow is predominantly diffusive. Panel (b) presents the corresponding $L^{1}$ residual norms for both the $L^{1}$- and $L^{2}$-trained models. The $L^{1}$-optimal Neural-POD exhibits improved accuracy in certain cases, particularly for specific reduced dimensions $r$ ($r=1$ and $r=3$) where it better captures localized steep gradients and nonlinear features. Because the $L^{1}$ norm penalizes large deviations less severely, it allows for sharper reconstructions in regions of strong nonlinearity, while the $L^{2}$ formulation maintains smoother but more globally averaged behavior. Overall, the comparison indicates that the relative performance of the two loss formulations depends on both the viscosity parameter and the dimension $r$.
(a) $L^{2}$ norm of the residuals for $L^{1}$- and $L^{2}$-optimal Neural-POD reconstructions.
(b) $L^{1}$ norm of the residuals for $L^{1}$- and $L^{2}$-optimal Neural-POD reconstructions.
Figure 9: Comparisons of $L^{1}$ and $L^{2}$ norms of the residuals for $L^{1}$- and $L^{2}$-optimal Neural-POD reconstructions.
3.1.2 Neural-POD for in-distribution and out-of-distribution viscosity values
In the following experiments, the Neural-POD model is trained under the same numerical setting as in Figure 5, using snapshot data generated at viscosity values $\nu$ sampled within the interval $[0.005,,0.01]$. This interval defines the training parameter domain for Neural-POD basis learning. Figure 10 shows both the in-distribution and out-of-distribution performance of the Neural-POD framework. Here, in-distribution refers to evaluating the learned Neural-POD model at viscosity values $\nu$ within the parameter interval used for training, while out-of-distribution refers to evaluating the model at $\nu$ values outside that interval. More specifically, Figure 10(a) and Figure 10(b) compare the first three basis modes from traditional POD with those produced by Neural-POD trained using the $L^{2}$ loss. Because traditional POD constructs modes only at the discrete viscosity levels contained in the snapshot data, it cannot directly produce basis functions for intermediate or out-of-range parameter values. Neural-POD, on the other hand, learns a continuous mapping from $\nu$ to the modal space, allowing it to generate basis functions for both unseen in-distribution and out-of-distribution parameter regimes. Notably, even for the fully unseen parameter value $\nu=4\times 10^{-3}$ (see Figure 10(b)), which lies outside the training interval $[0.005,,0.01]$ and represents a $20\%$ out-of-distribution shift below its lower bound, Neural-POD still generates modes that remain consistent with the corresponding POD modes that confirms its generalization capability.
Figure 10: Panel a: in-distribution and panel b: out-of-distribution of the first three POD modes: comparison between traditional POD with extra snapshots and Neural-POD for an unseen viscosity value $\nu$.
3.2 Navier-Stokes equations
$\nabla\cdot u=0\,$, (24)
where $u$ is the velocity, $p$ the pressure, and $Re$ the Reynolds number. The computational domain is a $2.2\times 0.41$ rectangular channel with a radius =0.05=0.05 cylinder, centered at $(0.2,0.2)$, see Figure 11. We prescribe no-slip boundary conditions on the walls and cylinder, and the following inflow and outflow profiles:
$u_{1}(0,y,t)$ $=u_{1}(2.2,y,t)=\frac{6}{0.41^{2}}y(0.41-y),$ (25)
$u_{2}(0,y,t)$ $=u_{2}(2.2,y,t)=0,$ (26)
where $u=\langle u_{1},u_{2}\rangle$.
(c) Geometry of two-dimensional flow past the cylinder.
(d) Sampling points for the flow past the cylinder test problem.
Figure 11: (a) Geometry and (b) sampling points of the two-dimensional flow past a cylinder problem.
The dataset used is generated by a high-fidelity finite element simulation of the unsteady Navier–Stokes equations at $Re=50$. The spatial discretization employs a Taylor–Hood element pair (quadratic velocity, linear pressure) on an unstructured triangular mesh that provides sufficient resolution of the boundary layers and wake structures. The resulting finite element discretization yields approximately $1.2\times 10^{4}$ velocity degrees of freedom (DOFs). Temporal integration is performed using a second-order backward differentiation formula (BDF2). A total of $1000$ velocity snapshots are recorded uniformly in time over the interval $[10,20]$. To make a consistent comparison, these snapshots are used to generate both the Proper Orthogonal Decomposition (POD) basis and to train the Neural–POD. Figure 12 illustrates the first two spatial modes identified by the classical POD and Neural–POD. The close agreement between the two confirms the consistency of Neural–POD with the traditional POD framework.
(a) First two POD modes
(b) First two Neural-POD modes
Figure 12: First two modes of POD and Neural-POD for two dimensional flow past the cylinder with $Re=50$.
3.3 Neural-POD-ROM
[…]
and $\nu=10^{-3}$. This test problem has been used in [mou2021data, wang2016approximate, koc2023uniform, reyes2025priori].
3.3.1 Snapshot Generation
The truth data are obtained using a linear finite element spatial discretization with mesh size $h=1/512$ and a Crank–Nicolson time discretization with timestep size $\Delta t=10^{-4}$. A total of $101$ snapshots are collected at equally spaced time instances over the interval $t\in[0,1]$ for constructing the POD and Neural-POD modes. More details are provided in Section 5.2.
3.3.2 Discussion of Numerical Results
$\nu$ Model $r=2$ $r=4$ $r=6$
0.002 POD–ROM 2.10e-01 1.66e-01 1.76e-01
Neural–POD–ROM ($L^{2}$) 2.11e-01 1.67e-01 1.77e-01
0.0025 POD–ROM No in distribution ability
Neural–POD–ROM ($L^{2}$) 9.61e-02 1.18e-01 1.35e-01
0.005 POD–ROM No out of distribution ability
Neural–POD–ROM ($L^{2}$) 1.91e-01 1.21e-01 8.46e-02
[…]
3.4 Neural-POD-DeepONet
We consider the one-dimensional Burgers’ equations (18)–(20) following the same configuration described in Section 3. The dataset is adopted from [lu2022comprehensive], where the initial conditions are sampled from a Gaussian random field characterized by a Riesz kernel, $\mu=\mathcal{R}(0,,625(-\Delta+25I)^{-2})$, with $\mu$ denoting the probability measure over the function space and $\Delta$ and $I$ representing the Laplacian and identity operators, respectively. Both input and output fields are discretized on a uniform grid with 128 spatial points.
In contrast to the method proposed in [lu2022comprehensive], which employs snapshots at $t=1$ to construct the POD basis, our Neural-POD framework learns the basis functions directly from the temporal evolution of the Burgers’ equation across multiple time steps. We employ four latent modes in the reduced representation and train the model on 128-point spatial grids, while evaluating its reconstruction and generalization performance on coarser 64-point grids. This setup enables assessment of both the model’s spatial resolution robustness and its capacity to generalize across varying discretizations and viscosity parameters. More details are presented in Appendix 5.3.
4 Conclusion and Future Work
Future research will focus on the following different directions: First, we will extend the Neural-POD framework to more challenging and high-dimensional scientific problems such the turbulent three-dimensional Navier–Stokes or Boussinesq equations, where nonlinear modal interactions and multi-scale energy cascades provide a demanding test [mou2023energy, mou2023combining, snyder2023data, zhang2022convergence, bright2013compressive, lin2009efficient]. Another promising direction is the integration of Neural-POD with physics-informed neural networks (PINNs) to enforce physical constraints during basis construction, thereby improving interpretability and extrapolation for physics-guided model reduction [cai2021physics, raissi2019physics, leung2022nh, lu2025fpinn, lu2025mopinnenkf]. Furthermore, coupling Neural-POD with data assimilation techniques such as Ensemble Kalman Filters (EnKF) or variational methods may enable real-time state estimation and model correction in dynamical systems [popov2021multifidelity, bilionis2013multi, chen2018conditional]. Also, in digital-twin applications, Neural-POD can be employed to build efficient and adaptive surrogate models that enable real-time monitoring, prediction, and control of complex physical systems [chen2025targeted, xiu2025computational, jones2020characterising]. Lastly, we plan to extend the framework to parameterize and learn geometrical variations that enables the community to contribute new mesh configurations and expand the applicability of Neural-POD to complex domains with heterogeneous geometries [shukla2024deep].

### v2

@@ Methods
The classical Proper Orthogonal Decomposition (POD) is a widely used technique for extracting low-dimensional structures from high-dimensional spatiotemporal data [berkooz1993proper, chatterjee2000introduction]. It is especially prominent in fluid dynamics, model reduction, and system identification due to its mathematical rigor, computational efficiency, and optimality properties in the $L^{2}$ norm.
@@
subject to the orthonormality constraint $(\phi_{i},\phi_{j})=\delta_{ij}$. The solution is obtained via singular value decomposition (SVD) of the snapshot matrix, and the resulting modes $\phi_{i}$} span an $r$-dimensional subspace that captures the most energetic components of the dataset [gubisch2017proper, kunisch2008proper, berkooz1992observations]. POD guarantees optimality in the $L^{2}$ sense: for any given rank $r$, the POD basis minimizes the projection error among all $r$-dimensional linear subspaces. This makes it a powerful tool for linear dimensionality reduction and the foundation for many reduced order models (ROMs), where the governing equations are projected onto the POD subspace using Galerkin or Petrov–Galerkin projection [girfoglio2021pod, iliescu2013variational, mou2020data, tsai2025time].
Despite its strengths, classical POD has several limitations. Because its basis is strictly linear, it often struggles to represent complex nonlinear features, sharp gradients, or discontinuities [volkwein2013proper, iliescu2014are]. Furthermore, the POD basis is highly dependent on the resolution and parameter regime of the snapshot data, limiting its generalizability [mifsud2016variable, roderick2014proper]. These limitations (see left panel of Figure 1(a)) motivate the development of more flexible, data-driven alternatives, such as the Neural-POD framework introduced in this work (see right panel of Figure 1(a)).
Neural-POD Framework
@@
Neural-POD in Galerkin Projection Based Reduced Order Models
@@
Traditional Projection-ROM
@@
where the tensors $C_{ijk}$ and $D_{ik}$ are now computed using the Neural-POD modes. The projection-based ROM using POD and Neural-POD is specifically illustrated in Figure 2(a).
@@
Neural-POD in Deep Operator Networks
@@
where $\phi_{k}(\cdot;\nu,\mathcal{X})$ denote the pretrained Neural-POD modes evaluated at viscosity $\nu$ and spatial grid $\mathcal{X}=\{x_{i}\}_{i=1}^{N}$, and $\phi_{0}$ is the corresponding mean field. The branch network outputs $b_{k}(v)$ represent the modal coefficients associated with these Neural-POD basis functions. The Neural-POD-DeepONet framework is illustrated in Figure 2(b).
@@ Acknowledgment
Data availability
Data used in this study are available at https://github.com/chhmou/Neural-POD . The full dataset is available upon reasonable request to the authors of the associated paper.
Code availability
The code used in this study, including scripts for data analysis in MATLAB (R2025b) and Python (v3.10), is available via GitHub at https://github.com/chhmou/Neural-POD. Python packages used for analysis include PyTorch, pandas, numpy, scikit-learn, matplotlib and statsmodels. Code annotations were generated using ChatGPT 4o (gpt-4o-2024-08-06).